OpenStudy (anonymous):

determine whether given series converges absolutely , converges conditionally or diverges?? ∑n=1∞ cosnΠ /n + 3

OpenStudy (sirm3d):

this? \[\large \sum_{n=1}^\infty \frac{\cos(n\pi)}{n+3}\] or this? \[\large \sum_{n=1}^\infty \cos\left(\frac{n\pi}{n+3}\right)\]

OpenStudy (anonymous):

upper one.

OpenStudy (sirm3d):

\[\cos n\pi=\pm 1\]

OpenStudy (goformit100):

@waterineyes

OpenStudy (anonymous):

@sirm3d right but what about Alternating series test..?

OpenStudy (sirm3d):

the series is alternating. given \[\sum (-1)^{n} a_n\] the series is (a) absolutely convergent if \[\sum a_n\] is convergent (b) conditionally convergent if the alternating series is convergent but the series of positive terms is divergent (c) divergent if the alternating series is divergent.

OpenStudy (sirm3d):

\[\sum \frac{\cos(n\pi)}{n+3}=\sum(-1)^{n}\frac{1}{n+3}\] take \[a_n=\frac{1}{n+3}\]

OpenStudy (anonymous):

thanks buddy problem solved :)

OpenStudy (sirm3d):

yw